# PatchRec: Multi-Grained Patching for Efficient LLM-based Sequential Recommendation

### v1

@@ Abstract.
Large Language Models for sequential recommendation (LLM4SR), which transform user-item interactions into language modeling, have shown promising results. However, due to the limitations of context window size and the computational costs associated with Large Language Models (LLMs), current approaches primarily truncate user history by only considering the textual information of items from the most recent interactions in the input prompt. This truncation fails to fully capture the long-term behavioral patterns of users. To address this, we propose a multi-grained patching framework — PatchRec. It compresses the textual tokens of an item title into a compact item patch, and further compresses multiple item patches into a denser session patch, with earlier interactions being compressed to a greater degree. The framework consists of two stages: (1) Patch Pre-training, which familiarizes LLMs with item-level compression patterns, and (2) Patch Fine-tuning, which teaches LLMs to model sequences at multiple granularities. Through this simple yet effective approach, empirical results demonstrate that PatchRec outperforms existing methods, achieving significant performance gains with fewer tokens fed to the LLM. Specifically, PatchRec shows up to a 32% improvement in HR@20 on the Goodreads dataset over uncompressed baseline, while using only 7% of the tokens. This multi-grained sequence modeling paradigm, with an adjustable compression ratio, enables LLMs to be efficiently deployed in real-world recommendation systems that handle extremely long user behavior sequences.
Sequential Recommendation, Large Language Model, Multi-Grained Compression
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 1. Introduction
(a) Truncating interaction sequences
(b) Performance of LLM4SR w.r.t. truncated sequence length
@@
Figure 1. (a) Most LLM4SR works consider the latest-$K$ interactions in the complete user interaction sequence as the truncated historical item sequence feed to LLMs. (b) Performance of LLM-based recommender on Movielens-1M at different numbers of items in the truncated interaction sequence. The number of tokens in historical item sequence ranges from 39 to 552, excluding the tokens of task description.
@@
Large Language Models (LLMs) (OpenAI, 2023; Google, 2023; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI, 2024) have demonstrated remarkable success across diverse domains, sparking a growing interest in their application to sequential recommendation (SR) (Wang et al., 2019a). A common paradigm for using LLMs as sequential recommender models (LLM4SR) (Geng et al., 2022; Yue et al., 2023; Zhang et al., 2023b; Bao et al., 2023b) involves two key steps: (1) formatting a user’s interaction history (i.e., a sequence of previously interacted items) into an input prompt suitable for LLMs, and (2) fine-tuning the LLMs to generate the target response (i.e., predicting the next item of interest). For the first step, most leading methods (Bao et al., 2023b; Liao et al., 2024b, a; Yang et al., 2023b) truncate each interaction history, retaining only the most recent $K$ items [^1] K, with common values being 5, 10, or 20, is typically much smaller than the total length of complete interaction history.. This truncation integrates a much shorter interaction sequence into the input prompt, primarily due to the context window size limitations and computational demand of LLMs. In such a truncated interaction sequence, each item is represented using the textual metadata (e.g., movie titles like “Gone with the Wind”) (Bao et al., 2023b, a; Liao et al., 2024a), as shown in Figure 1(a). While effective for capturing short-term preferences, this approach directly ignores interactions prior to the most recent $K$ items, making it challenging to capture the long-term interests encoded in extended item sequences. Modeling long sequences remains largely unexplored in LLM4SR, but the primary focus of this paper.
To investigate the impact of sequence length, we conduct preliminary experiments by supervised fine-tuning (SFT) Llama-3.2-1B-Instruct (Dubey et al., 2024) with different truncated sequence length on the Movielens dataset, as depicted in Figure 1(a). The results, illustrated in Figure 1(b), yield two key observations: (1) Performance improves with truncated sequence length: As the number of items in the truncated interaction sequence increases, recommendation accuracy — HR@20, improves substantially, showcasing the value of modeling longer interaction histories. For instance, earlier interactions with animated films contribute to accurately predicting subsequent items like “Alice Wonderland”. (2) Performance plateaus beyond a threshold: However, beyond a certain sequence length (e.g., 60 items), performance levels off with opaque increase. This plateau effect may result from the diminishing relevance of older interactions to a user’s current preferences and the potential introduction of noise (Zhou et al., 2018; Chen et al., 2022). Take Figure 1(a) as an example again, where older items like “Pulp Fiction” may serve as irrelevant or noisy signals in the user‘s interaction history.
These observations motivate us to consider the following challenges of modeling long sequences in LLM4SR:
  • Constraints of LLMs: The length of the item sequence significantly influences the input prompt length, which is constrained by the context window length of LLMs (Su et al., 2024) and the associated computational costs. How can we design approaches to efficiently model longer item sequences within these limitations to mine long-term user interest?
  • Temporal significance of interactions: Current LLM4SR methods do not explicitly regard the temporal distinctions of relevance (Ying et al., 2018), treating all items in a user’s interaction sequence as equally important (Liao et al., 2024b; Bao et al., 2023a). How can we incorporate mechanisms to capture the varying importance of interactions over time, thereby aligning more closely with evolving user preferences?
Figure 2. Performance-efficiency trade-off for vanilla LLM4SR method, PatchRec-I, and PatchRec-S on MovieLens-1M. PatchRec variants demonstrate improvements in both metrics compared to the baseline. The shaded region highlights the area above the performance-efficiency trade-off curve, representing desirable configurations.
@@
To tackle these challenges, we introduce a hierarchical compression approach, as depicted in Figure 3. First, the textual tokens of item titles are compressed into compact item patches. These items are then organized into sessions, with the corresponding item patches further compressed into session patches, capturing higher-level patterns in user behavior. This compression allows for the accommodation of a significantly larger number of items within a fixed context window. Additionally, to account for the temporal significance of user interactions, we integrate multi-grained session patches, item patches, and raw textual tokens, enabling more compact and dynamic modeling of user interaction histories.
Building on this idea, we propose PatchRec — a new patching framework for LLM4SR. As illustrated in Figure 4, PatchRec operates in the following two stages:
@@
Extensive experiments are conducted on three benchmark datasets (i.e., MovieLens-1M (Harper and Konstan, 2016), Goodreads (Wan and McAuley, 2018), and MovieLens-100K (Harper and Konstan, 2016)), revealing the following key findings: (1) For the same truncated interaction sequence, PatchRec uses only 7.34% LLM input tokens compared to uncompressed SFT, while simultaneously improving HR@20 by up to 32% on the Goodreads dataset. (2) For the same computational cost, PatchRec models 3.44 times more user behaviors than uncompressed SFT, resulting in a substantial performance improvement of up to 13% on the MovieLens-1M dataset.
@@
    We propose a multi-grained (i.e., both item- and session-level) compression framework — PatchRec, modeling users’ long-term historical behaviors in LLM4SR.
@@
2. Related Work
In this section, we provide a brief overview of existing works on LLM-based sequential recommendation (LLM4SR) and discuss user sequence modeling approaches in sequential recommendation.
2.1. LLM for Sequential Recommendation
Sequential recommendation involves predicting the next item a user is likely to interact with, based on their past engagements (Fang et al., 2020; Wang et al., 2019b). With LLMs demonstrating vast world knowledge and impressive reasoning capabilities across various domains (OpenAI, 2023; Google, 2023; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI, 2024; Shao et al., 2024), their potential for sequential recommendation tasks has garnered significant attention from researchers (Hou et al., 2024; Yue et al., 2023). Existing work on LLM4SR can be roughly classified into two categories: (1) LLMs as enhancers: In this approach, LLMs augment traditional recommender models by processing textual features associated with users and items (Yuan et al., 2023; Hou et al., 2022), and (2) LLMs as recommenders: LLMs directly generate recommendations for users’ next interactions (Gao et al., 2023; Bao et al., 2023b; Rajput et al., 2023).
The primary focus of this paper is on the latter — LLMs as recommenders — where the core challenge is converting user-item interaction sequences into a format suitable for language modeling task. Current approaches in LLM4SR can be broadly categorized into two types: (1) Text-based item representations: Items are represented using texts, such as item titles (Bao et al., 2023b, a), descriptions (Li et al., 2023), or index numbers (Geng et al., 2022), which are readily digestible for LLMs, and (2) Multi-modal token incorporation: Items are represented by tokens derived from other modalities (Liao et al., 2024b; Zhang et al., 2023a), such as embeddings projected from item representations learned by traditional recommender models (Tang and Wang, 2018; Hidasi et al., 2016a; Kang and McAuley, 2018).
In both aforementioned categories, each item is typically represented by multiple tokens. Due to the quadratic computational complexity $O(n^{2})$ of self-attention mechanism in LLMs w.r.t. sequence length  [^2] The computational complexity of vanilla attention mechanisms is $O(n^{2})$. Recent studies proposed methods to reduce it to approximately $O(n)$, using multiple tokens to represent one item in a user’s interaction sequence becomes computationally demanding. This motivates us to explore more efficient item and even sequence representations, aiming to improve both the performance and efficiency of LLM4SR methods.
2.2. User Sequence Modeling
Sequential recommendation interprets users’ interaction sequences as reflections of their underlying interests. Traditional sequential recommender models leverage neural networks to capture these patterns, employing architectures such as Recurrent Neural Networks (RNNs) (Hidasi et al., 2016a; Tan et al., 2016; Quadrana et al., 2017; Beutel et al., 2018; Hidasi et al., 2016b), Convolutional Neural Networks (CNNs) (Tang and Wang, 2018; Tuan and Phuong, 2017; Yuan et al., 2019), and attention mechanisms (Kang and McAuley, 2018; Sun et al., 2019). These models primarily focus on users’ latest-$K$ interactions, thus emphasizing short-term interests.
To incorporate information from richer historical behaviors, several studies have extended these approaches to model long-term interests. Methods include introducing static user representations (Zhang et al., 2018), leveraging two-layer hierarchical attention networks (Ying et al., 2018), periodically updating user memory (Ren et al., 2019), and employing multi-channel memory networks (Pi et al., 2019). To efficiently model long-term user behaviors, some works adopt linear attention mechanism (Liu et al., 2023) and mamba (Liu et al., 2024) in the architecture. Additionally, other works retrieve items from a user’s interaction history based on their relevance to the target item (Pi et al., 2020; Qin et al., 2022; Chen et al., 2021; Cao et al., 2022).
The integration of LLMs into recommendation systems introduces a unique challenge — further constraining the number of users’ historical behaviors that can be modeled due to the context window length of LLMs and the associated computational resource limitations. Despite this challenge, few studies have explored the modeling of long sequences in LLM4SR. One notable work (Lin et al., 2024) tackles the challenge by retrieving the top-$K$ interacted items from a user’s history based on semantic relevance to the target item. However, this retrieval-based approach is specifically designed for the click-through rate (CTR) prediction task and is not directly applicable to the next-item prediction task in sequential recommendation, which is the primary focus of this paper.
3. Method: PatchRec
Figure 3. Hierarchical compression. The textual tokens of an item title are aggregated into a compact item patch. Then several adjacent item patches are further compressed into a denser session patch.
We propose PatchRec, a simple yet effective framework designed to model longer historical item sequences in LLM-based sequential recommendation (LLM4SR) through hierarchical compression. This section begins with the task formation of LLM4SR, followed by an introduction to our hierarchical compression method. Finally, we detail the two-stage training process that underpins our framework.
3.1. Task Formation
Current LLMs, parameterized by $\Theta$, are typically trained using an autoregressive modeling approach, where the model generates each token $\mathbf{x}_{t}$ conditioned on the previous ones $\mathbf{x}_{<t}$ of the sequence. Formally, this is expressed as:
(1) $P(X;\Theta)=\prod_{t=1}^{T}P(\mathbf{x}_{t}|\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{t-1};\Theta).$
In the context of sequential recommendation, LLM4SR utilizes LLMs as sequential recommenders to generate responses, i.e., predicting the next item that aligns with user preferences, based on an input prompt describing their historical interaction sequence. This task can be formalized as language modeling:
@@ 3.1. Task Formation
where $X$ represents the LLM input sequence of $N$ tokens, $Y$ denotes the output sequence, and $\Theta$ are the model parameters. In this formulation, $X$ typically consists of a truncated interaction sequence of $K$ items accommodated in a prompt template. Each item title in $X$ is often represented by several textual tokens $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the hidden size of the LLM. The sequence $Y$ is intended to represent the predicted title of the next item.
3.2. Hierarchical Compression
To incorporate more interactions within truncated sequences, given the limitation of LLMs’ context window length and computation demand, we hierarchically compress the representation of historical item sequences, as shown in Figure 3.
3.2.1. Item Patch Compression.
In the context of LLM4SR, the textual tokens of each item title naturally form a unit that semantically represents the item. To compress these textual tokens into an item patch, we propose a simple yet effective method — computing the average of their token embeddings as follows:
@@ 3.2. Hierarchical Compression
where $\mathbf{z}^{i}_{j}\in\mathbb{R}^{d}$ denotes the item patch for the $j$-th item, $\mathcal{T}_{j}$ is the set of textual tokens associated with this item’s title, and $\mathbf{x}_{t}$ is the textual embedding of token $t$. As shown in Figure 3, the textual tokens of a movie title (e.g., “The Bridges of Madison Country”) are compressed into one compact item patch. This compression reduces the token-level granularity into a semantic-level representation of the item, effectively compressing the input space.
3.2.2. Session Patch Compression.
Since users often exhibit consistent interests over a period of time, consecutive interactions can be further compressed. To accomplish this, we partition a user’s historical item sequence into sessions, where each session groups together a predefined number of $L$ adjacent items  [^3] Alternative partitioning methods can also be employed, such as grouping items interacted with during a specific time window.. The item patches within each session are subsequently compressed into a session patch by averaging the embeddings of the item patches as follows:
@@
3.3. Two-Stage Training
Although the compressed item/session patches $\mathbf{z}^{i}$/$\mathbf{z}^{s}$ maintain the same dimensionality as the original textual token, LLMs do not inherently understand them. To address this, we propose a two-stage training process, as illustrated in Figure 4. In the first stage, Patch Pre-training, we train the LLM to learn the compression patterns from textual tokens to item patches by comparing item patches with their corresponding original textual tokens. In the second stage, Patch Fine-tuning, we adapt the LLM to the final multi-grained and time-aware compressed format.
@@ 3.3.2. Patch Fine-tuning
To incorporate the temporal decay of user interests in a sequence, we propose two fine-tuning strategies with different compression granularities: Patch Fine-tuning on Items (PFT-I) and Patch Fine-tuning on Sessions (PFT-S). As shown in Figue 4, they are designed to accommodate varying computational resource constraints.
@@
4. Experiments
In this section, we present a comprehensive comparison of PatchRec against various baseline methods, including traditional recommender models, efficient long-term recommender models, and LLM-based recommenders, across three real-world datasets: MovieLens-1M, Goodreads, and MovieLens-100K. Our objective is to demonstrate the effectiveness and efficiency of PatchRec under different levels and granularities of compression. Additionally, we analyze the mechanisms through which our approach achieves its superior performance. Furthermore, we analyze the performance of PatchRec in user groups with more long-term behaviors.
4.1. Experimental Settings
Table 1. Statistics of datasets. Int.: Interactions. TT.: Title tokens.
Dataset # User # Item # Int. # Max Int. # Avg. Int. # Avg. TT.
@@ 4.1. Experimental Settings
4.1.1. Datasets
We conduct extensive experiments on three real-world benchmark datasets: MovieLens-1M, Goodreads, and MovieLens-100K. The MovieLens datasets are widely used benchmarks, sourced from a movie recommendation platform  [^5] https://movielens.org/, while the Goodreads dataset is derived from the largest online community for readers and book recommendations  [^6] https://www.goodreads.com/.
4.1.2. Implementations
We adopt Llama-3.2-1B-Instruct (Dubey et al., 2024) as the backbone LLM for our framework, fine-tuning all its parameters. Each training stage consists of one epoch and requires approximately 12 NVIDIA A800 GPU hours. To balance computational efficiency and resource allocation, we use a training batch size of 64 for MovieLens datasets and 32 for Goodreads, and a test batch size of 16 for all three datasets. The training employs a cosine learning rate scheduler with a warmup ratio of 0.05, and the learning rates are set to 8e-6,5e-5,1e-5 for MovieLens-1M, Goodreads, and MovieLens-100K, respectively. For each interaction, the latest $K$ historical interactions are selected as the truncated user sequence. All experiments access item titles that appear within their respective datasets to ensure fair and consistent evaluation.
@@ 4.1.2. Implementations
4.1.3. Evaluation
To evaluate recommendation performance, we use two key metrics: HitRatio@10/20 and NDCG@10/20. Additionally, to assess efficiency, we introduce the Compression Ratio, defined as the ratio of the number of tokens in the historical item sequence before compression to the number after compression. A higher compression ratio indicates greater efficiency.
4.2. Main Performance Comparison
4.2.1. Baselines
We denote the models trained using patch pre-training and PFT-I as PatchRec-I, and those trained with patch pre-training and PFT-S as PatchRec-S. Our method is compared against a diverse set of baselines, including traditional recommender models, efficient long-term sequential recommender models, and LLM-based recommenders.
  • Traditional Recommender Models. We select three sequential recommender models as representatives of different mechanisms for capturing user behavior patterns: GRU4Rec(Hidasi et al., 2016a) (RNN-based), Caser(Tang and Wang, 2018) (CNN-based) and SASRec(Kang and McAuley, 2018) (attention-based)  [^7] These models are implemented following (Yang et al., 2023a)..
  • Efficient Long-term Recommender Models. LinRec (Liu et al., 2023) and Mamba4Rec (Liu et al., 2024) are chosen as representatives of efficient long-term sequential recommendation methods. LinRec leverages linear attention mechanisms, while Mamba4Rec employs Mamba blocks (Gu and Dao, 2023) for efficient modeling of long-term user behaviors.
  • LLM-based Recommenders. We include three representative methods that incorporate LLMs in sequential recommendation: (1) MoRec  [^8] We adopt BERT as the encoder for items’ textual metadata and SASRec as the recommender backbone. (Yuan et al., 2023) enhances traditional recommenders by incorporating textual features encoded by LLMs. (2) Vanilla LLM4SR (Vanilla) is adapted from leading methods such as TALLRec (Bao et al., 2023b) and BIGRec (Bao et al., 2023a) by applying SFT on truncated interaction sequences without compression, while maintaining evaluation settings consistent with PatchRec. It serves as the most important control group for PatchRec, with the only difference being whether the historical item sequence is compressed.

### v2

# Multi-Grained Patch Training for Efficient LLM-based Recommendation

@@ Abstract.
Large Language Models (LLMs) have emerged as a new paradigm for recommendation by converting interacted item history into language modeling. However, constrained by the limited context length of LLMs, existing approaches have to truncate item history in the prompt, focusing only on recent interactions and sacrificing the ability to model long-term history. To enable LLMs to model long histories, we pursue a concise embedding representation for items and sessions. In the LLM embedding space, we construct an item’s embedding by aggregating its textual token embeddings; similarly, we construct a session’s embedding by aggregating its item embeddings. While efficient, this way poses two challenges since it ignores the temporal significance of user interactions and LLMs do not natively interpret our custom embeddings. To overcome these, we propose PatchRec, a multi-grained patch training method consisting of two stages: (1) Patch Pre-training, which familiarizes LLMs with aggregated embeddings – patches, and (2) Patch Fine-tuning, which enables LLMs to capture time-aware significance in interaction history. Extensive experiments show that PatchRec effectively models longer behavior histories with improved efficiency. This work facilitates the practical use of LLMs for modeling long behavior histories. Codes are available at https://github.com/ljy0ustc/PatchRec.
Sequential Recommendation; Large Language Model; Multi-Grained Compression
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730042isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. Most LLM4SR works consider the latest-$K$ interactions in the complete user interaction sequence as the truncated historical item sequence fed to LLMs.
@@
Large Language Models (LLMs) (OpenAI, 2023; Google, 2023; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI, 2024) have demonstrated remarkable success across diverse domains, sparking a growing interest in their application to recommendation. A common paradigm for using LLMs as recommender models (LLM4Rec) (Geng et al., 2022; Yue et al., 2023; Zhang et al., 2023; Bao et al., 2023b) involves two key steps: (1) formatting a user’s interaction history (i.e., a sequence of previously interacted items) into an input prompt suitable for LLMs, and (2) fine-tuning the LLMs to generate the target response (i.e., predicting the next item of interest). For the first step, most leading methods (Bao et al., 2023b; Liao et al., 2024b, a; Yang et al., 2023b) truncate each interaction history, retaining only the most recent $K$ items [^1] K, with common values being 5, 10, or 20, is typically much smaller than the total length of complete interaction history., with each item represented by the textual metadata (e.g., movie titles like “Gone with the Wind”) (Bao et al., 2023b, a; Liao et al., 2024a), as shown in Figure 1. This truncation integrates a much shorter interaction sequence into the input prompt, primarily due to LLMs’ context window size limitations and the associated computational costs. While effective for capturing short-term preferences, this approach abandons the interactions prior to the selected items, failing to capture the long-term interests encoded in extended item sequences.
@@
To efficiently model long item sequences in LLM4Rec, we introduce a hierarchical representation approach, as depicted in Figure 4. In the embedding space, the textual tokens of item titles are firstly compressed into compact item patches. These items are then organized into sessions, with the corresponding item patches further compressed into session patches, capturing higher-level patterns in user behavior. This compressed representation allows for the accommodation of a significantly larger number of items within a fixed context window. However, the following two challenges remain:
  • Temporal significance of interactions: Such compressed representations do not explicitly regard the temporal distinctions of relevance (Ying et al., 2018), treating all items in a user’s interaction sequence as equally important (Liao et al., 2024b; Bao et al., 2023a). How can we incorporate mechanisms to capture the varying importance of interactions over time, thereby aligning more closely with evolving user preferences?
  • Comprehension of compression patterns: While compressed item and session patches retain the same dimensionality as original textual tokens, LLMs lack inherent understanding of them. How can we design training approaches to enable LLMs to interpret these compressed patterns effectively?
To tackle these challenges, we propose PatchRec, a multi-grained patch training framework that unifies: (1) session patches for high-level behavioral patterns, (2) item patches for fine-grained preference modeling, and (3) raw textual tokens for semantic grounding. This framework enables compact yet expressive representation of user interaction history, while maintaining dynamic adaptability to varying sequence lengths. As illustrated in Figure 5, PatchRec operates in the following two stages:
@@
Figure 2. Performance-efficiency trade-off for TALLRec, Pure-Item/Session, and PatchRec-I/S on MovieLens-1M. PatchRec variants demonstrate improvements in both metrics compared to the baseline. The shaded region highlights the area above the performance-efficiency trade-off curve, representing desirable configurations.
Extensive experiments are conducted on datasets and yield the following key findings: (1) For the same truncated interaction sequence, PatchRec uses only 7.34% LLM input tokens compared to TALLRec, while simultaneously improving HR@20 by up to 32% on the Goodreads dataset. (2) For the same computational cost, PatchRec models 3.44 times more user behaviors than TALLRec, resulting in a substantial performance improvement of up to 13% on the MovieLens-1M dataset.
@@
    We propose a multi-grained (i.e., both item- and session-level) patch training framework — PatchRec, modeling users’ long-term historical behaviors in LLM4Rec.
@@
2. Preliminary
Given an input prompt describing a user’s chronological interaction history with items, an LLM-based recommender predicts the next item that best matches this user’s preference. Building upon a foundational work in LLM-based recommendation, TALLRec (Bao et al., 2023b), we formalize this baseline as supervised finetuning an LLM parameterized by $\Theta$ to maximize the autoregressive likelihood:
@@ 2. Preliminary
where $X=x_{1:N}$ is the LLM input sequence of $N$ tokens representing the user interaction history, and $Y=y_{1:T}$ is the output sequence denoting the next item. The input sequence $X$ encodes a truncated interaction sequence of $K$ items using a structured prompt template. Each item is represented through its title, which is subsequently tokenized into several tokens $\mathbf{x}_{i}\in\mathbb{R}^{d}$, where $d$ is the hidden size of the LLM. To adapt TALLRec from click-through rate prediction (answering YES/NO given a target item) to next-item recommendation, we let the LLM generate the next item’s title $Y$ token-by-token through autoregressive decoding, employing constrained beam search over a token-level prefix tree constructed from all candidate item titles. This ensures that generated sequences correspond to valid items.
Figure 3. Performance of LLM-based recommender on Movielens-1M at different numbers of items in the truncated interaction sequence. The number of tokens in the historical item sequence ranges from 39 to 552, excluding the tokens of task description.
To investigate the impact of sequence length, we scale up $K$ in preliminary experiments. Figure 3 reveals a key trend: performance improves with truncated sequence length. As the number of items in the truncated interaction sequence increases, recommendation accuracy — HR@20, improves substantially, suggesting that richer user histories enable more precise modeling of preferences. For instance, earlier interactions with animated films contribute to accurately predicting subsequent items like “Alice in Wonderland”.
Despite the benefits of modeling long sequences, this baseline approach faces a critical limitation: excessive inference costs due to rapidly growing LLM prompt lengths. The length of the item sequence significantly influences the input prompt length, which is constrained by the context window length of LLMs (Su et al., 2024) and the associated computational costs.
3. Method
In this section, we first present the cornerstone of our approach — the hierarchical patching strategy — and then describe the two-stage PatchRec framework.
3.1. Hierarchical Patching
To incorporate more interactions within constraints of LLM prompt length overhead, we hierarchically compress the representation of historical item sequences, as shown in Figure 4.
Figure 4. Hierarchical compression. The textual tokens of an item title are aggregated into a compact item patch. Then, several adjacent item patches are further compressed into a denser session patch.
3.1.1. Item Patch Construction.
In the context of LLM4Rec, the textual tokens of each item title naturally form a semantically coherent unit. To compress these tokens into a single token representation – termed an item patch – we adopt a simple yet effective method of average pooling. Specifically, we compute the mean of the token embeddings corresponding to the item’s title as follows:
@@ 3.1. Hierarchical Patching
where $\mathbf{z}^{i}_{j}\in\mathbb{R}^{d}$ is the resulting item patch for the $j$-th item, $\mathcal{T}_{j}$ denotes the set of tokens in the item’s title, and $\mathbf{x}_{t}$ is the embedding of token $t$. This average pooling approach compresses token-level information into a compact item-level representation. As shown in Figure 4, a movie title such as “The Bridges of Madison Country” is encoded into a single item patch. This reduces the input length while maintaining the core content of the item description.
3.1.2. Session Patch Construction.
Since users often exhibit consistent interests over a period of time, consecutive interactions can be further compressed. To accomplish this, we partition a user’s historical item sequence into sessions, where each session is formed by grouping a fixed number $L$ of consecutive items  [^2] Alternative partitioning methods can also be employed, such as grouping items interacted with during a specific time window.. We then obtain a session patch by averaging the embeddings of the item patches within each session:
@@
3.1.3. Challenges of Patch Training
Based on our proposed item and session patch compression methods, we first examine two intuitive adaptations: substituting all text tokens in a sequence with item or session patches only, named Pure-Item and Pure-Session. While effectively reducing LLM input tokens, they ignore the temporal significance of user interactions and the LLM’s inherent incomprehension of compression patterns, suggesting clear opportunities for improvement. As shown in Figure 2, these approaches achieve suboptimal performance-efficiency trade-offs, remaining below the Pareto frontier in the upper-right optimization space.
3.2. PatchRec
To accommodate both challenges of time-aware sequence modeling and compression pattern understanding, we propose PatchRec, a simple yet effective framework featuring two-stage training, as illustrated in Figure 5. To model the varying relevance of earlier versus later interacted items to a user’s current interests, we employ a multi-grained patching strategy. Specifically, earlier interactions are compressed at higher levels (e.g., as item and even session-level patches) for capturing long-term preferences, while most recent interactions retain finer granularity (e.g., textual tokens) for preserving semantic grounding. Since LLMs do not inherently understand the compression pattern from textual tokens to item patches, in the first stage, Patch Pre-training, we train the LLM to learn it by comparing item patches with their corresponding original textual tokens. In the second stage, Patch Fine-tuning, we further adapt the LLM to the final multi-grained compressed format.
@@ 3.2.2. Patch Fine-tuning
To incorporate the temporal decay of user interests in a sequence, we propose two fine-tuning strategies with different compression granularities: Patch Fine-tuning on Items (PFT-I) and Patch Fine-tuning on Sessions (PFT-S). As shown in Figue 5, they are designed to accommodate varying computational resource constraints.
@@
Table 1. Statistics of datasets. Int.: Interactions. SL.: Items per user sequence. TT.: Title tokens per item.
Dataset # User # Item # Int. # Max SL. # Avg. SL. # Avg. TT.
@@
4. Experiments
In this section, we present a comprehensive comparison of PatchRec against various baseline methods across real-world datasets, demonstrating the effectiveness and efficiency of PatchRec under different levels and granularities of compression. Additionally, we analyze the mechanisms through which our approach achieves its superior performance and its performance in user groups with more long-term behaviors.
4.1. Experimental Settings
Datasets. We conduct extensive experiments on three real-world benchmark datasets: MovieLens-1M, Goodreads, and MovieLens-100K. The MovieLens datasets are widely used benchmarks, sourced from a movie recommendation platform  [^4] https://movielens.org/, while the Goodreads dataset is derived from the largest online community for readers and book recommendations  [^5] https://www.goodreads.com/.